# Group cohesion and passive dynamics of a pair of inertial swimmers with three-dimensional hydrodynamic interactions

### v1

@@ IV Conclusion
The two-dimensional swimmers largely have the same planar dynamics as the three-dimensional swimmers. However, our three-dimensional model enables the study of non-planar dynamics and three-dimensional schools, which we will pursue in future work.
@@ Appendix A Phase diagrams for three-dimensional model
This appendix contains the phase diagrams for the planar dynamics of a pair of three-dimensional source-sink swimmers with $\Delta\theta\neq 0$. Symbols and colors are as in Figure 10.
Figure A.1: Phase diagrams for $|\Delta\theta|=\frac{\pi}{12}$ for three-dimensional swimmers.
Figure A.2: Phase diagrams for $|\Delta\theta|=\frac{\pi}{6}$ for three-dimensional swimmers.
Figure A.3: Phase diagrams for $|\Delta\theta|=\frac{\pi}{4}$ for three-dimensional swimmers.
Figure A.4: Phase diagrams for $|\Delta\theta|=\frac{\pi}{3}$ for three-dimensional swimmers.
Figure A.5: Phase diagrams for $|\Delta\theta|=\frac{5\pi}{12}$ for three-dimensional swimmers.
Figure A.6: Phase diagrams for $|\Delta\theta|=\frac{\pi}{2}$ for three-dimensional swimmers.
Figure A.7: Phase diagrams for $|\Delta\theta|=\frac{7\pi}{12}$ for three-dimensional swimmers.
Figure A.8: Phase diagrams for $|\Delta\theta|=\frac{2\pi}{3}$ for three-dimensional swimmers.
Figure A.9: Phase diagrams for $|\Delta\theta|=\frac{3\pi}{4}$ for three-dimensional swimmers.
Figure A.10: Phase diagrams for $|\Delta\theta|=\frac{5\pi}{6}$ for three-dimensional swimmers.
Figure A.11: Phase diagrams for $|\Delta\theta|=\frac{11\pi}{12}$ for three-dimensional swimmers.
Figure A.12: Phase diagrams for $|\Delta\theta|=\pi$ for three-dimensional swimmers.
@@ Appendix B Phase diagrams for two-dimensional model
This appendix contains the phase diagrams for the planar dynamics of a pair of two-dimensional source-sink swimmers with $\Delta\theta\neq 0$. Symbols and colors are as in Figure 14.
Figure B.1: Phase diagrams for $|\Delta\theta|=\frac{\pi}{12}$ for two-dimensional swimmers.
Figure B.2: Phase diagrams for $|\Delta\theta|=\frac{\pi}{6}$ for two-dimensional swimmers.
Figure B.3: Phase diagrams for $|\Delta\theta|=\frac{\pi}{4}$ for two-dimensional swimmers.
Figure B.4: Phase diagrams for $|\Delta\theta|=\frac{\pi}{3}$ for two-dimensional swimmers.
Figure B.5: Phase diagrams for $|\Delta\theta|=\frac{5\pi}{12}$ for two-dimensional swimmers.
Figure B.6: Phase diagrams for $|\Delta\theta|=\frac{\pi}{2}$ for two-dimensional swimmers.
Figure B.7: Phase diagrams for $|\Delta\theta|=\frac{7\pi}{12}$ for two-dimensional swimmers.
Figure B.8: Phase diagrams for $|\Delta\theta|=\frac{2\pi}{3}$ for two-dimensional swimmers.
Figure B.9: Phase diagrams for $|\Delta\theta|=\frac{3\pi}{4}$ for two-dimensional swimmers.
Figure B.10: Phase diagrams for $|\Delta\theta|=\frac{5\pi}{6}$ for two-dimensional swimmers.
Figure B.11: Phase diagrams for $|\Delta\theta|=\frac{11\pi}{12}$ for two-dimensional swimmers.
Figure B.12: Phase diagrams for $|\Delta\theta|=\pi$ for two-dimensional swimmers.

### v2

@@ IV Conclusion
The two-dimensional swimmers largely have the same planar dynamics as the three-dimensional swimmers. However, our three-dimensional model enables the study of non-planar dynamics and three-dimensional schools. The non-planar dynamics are of the same character as the planar dynamics, although we observe the emergence of multiscale dynamics in the fully three-dimensional problem. Despite the potential for greater misalignment between swimmers, we observe configurations that lead to a passively cohesive group. A detailed investigation of the non-planar problem will be the subject of future work.
@@
This work was supported by the University of Houston Grants to Enhance and Advance Research Program, 000189684.
@@ Appendix A Phase diagrams for three-dimensional model
The phase diagrams for the planar dynamics of a pair of three-dimensional source-sink swimmers with $\Delta\theta\neq 0$ are shown in Figure A.1 in increments of $\pi/12$. Symbols and colors are as in Figure 9. Outside of the regions shown, the swimmers diverge from each other; that is, those regions of the phase diagrams are filled with red diamonds. The phase diagrams for $\Delta\theta<0$ (not shown) are the same as for $\Delta\theta>0$, but reflected across the $\Delta x=0$ axis.
Figure A.1: Phase diagrams for three-dimensional swimmers for $\Delta\theta>0$.
@@ Appendix B Phase diagrams for two-dimensional model
The phase diagrams for the planar dynamics of a pair of two-dimensional source-sink swimmers with $\Delta\theta\neq 0$ are shown in Figure B.1 in increments of $\pi/12$. Symbols and colors are as in Figure 13. Outside of the regions shown, the swimmers diverge from each other; that is, those regions of the phase diagrams are filled with red diamonds. The phase diagrams for $\Delta\theta<0$ (not shown) are the same as for $\Delta\theta>0$, but reflected across the $\Delta x=0$ axis.
Figure B.1: Phase diagrams for two-dimensional swimmers for $\Delta\theta>0$.